OpenStudy (anonymous):

how do you solve by substitution and graphing with this equation: eq 1: y= 1/2x + 4 eq 2: y= -x - 5

OpenStudy (anonymous):

Take equation (2). Substitute the value of y in equation (1). You will get an equation in x. Solve for x. Then, solve for y using equation (2).

OpenStudy (anonymous):

To solve by graphing, you should find the x and y intercepts for both equations, graph the lines and find the point of intersection.

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Im confused still. When i add the opposite to 1/2x to get x by itself, i get so confused with all the fractions

OpenStudy (anonymous):

substituting equation (2) in equation (1), you get: -x - 5 = 1/2 x + 4 So, rearranging, you have: 1/2 x + x = -5-4 3/2 x = -9 x = -6

OpenStudy (anonymous):

y = 1/2 X + 4 , y = -X + 5 , 1/2 X + 4 = -X +5 , 3/2 X = -9 , X = 2 ( -9) / 3 = -6 ---> y = 1